# MagicDriveDiT: High-Resolution Long Video Generation for Autonomous Driving with Adaptive Control

### v1

@@ E More Comparison among VAEs
Table II: VAE Comparison for Street Views. CogVAE [43] and Open-Sora [47] (1.2) are 3D VAEs; SD VAE [31] is 2D VAE, which is also widely adopted by previous street view generation (e.g., [11]). Results are PSNRs calculated through videos from the nuScenes validation set. MagicDriveDiT adopts CogVAE.
F Reason for Using CogVAE without the Pre-trained Diffusion Model
@@
We conducted a comparative analysis of VAEs, as detailed in Section 5.3 and Appendix E, and found CogVAE [43] to perform well. Given that the VAE determines the upper limit of generation quality, we opted to use CogVAE for video encoding. Notably, CogVideoX [43] employs a novel DiT structure, where each layer’s latent space integrates both video and text condition information. This approach may complicate the design of geometry-related conditions. Furthermore, CogVideoX was not trained in a driving scenario. To eliminate these potential confounding factors, we decided to train the diffusion model from scratch using CogVAE. This strategy allows us to move beyond the constraints of pre-trained models, enabling more flexible modifications to the model architecture to achieve multi-view consistency and spatiotemporal encoding of geometry conditions.
@@
G Single Inference v.s. Rollout Inference
(a) Generation from Vista. It takes the first frame as input and generates the following (only support the front view).
@@
Figure II: Comparison between Rollout for Long Videos (Vista [12]) and Single Inference (our MagicDriveDiT). Although rollout can handle long videos, the quality is significantly degraded. In contrast, our extrapolation maintains high quality in long video generation.
@@
H More Visualization
As said in the “Note”, MagicDriveDiT is capable of generating $6\times 848\times 1600\times 241$ videos (20s at 12 fps). We include more generated samples in Figure III-IV. Please see the videos on our project page.
Figure III: We show some frames from the generated $6\times 848\times 1600\times 241$ videos with the same scene configuration (i.e., boxes, maps, cameras, and ego trajectory) but under different weather conditions. Conditions are from the nuScenes validation set.
Figure IV: We show some frames from the generated $6\times 848\times 1600\times 241$ videos with the same scene configuration (i.e., boxes, maps, cameras, and ego trajectory) but under different time-of-day conditions. Conditions are from the nuScenes validation set.

### v2

@@ G More Comparison among VAEs
Table III: VAE Comparison for Street Views. CogVAE [42] and Open-Sora [46] (1.2) are 3D VAEs; SD VAE [30] is 2D VAE, which is also widely adopted by previous street view generation (e.g., [11]). Results are PSNRs calculated through videos from the nuScenes validation set. MagicDriveDiT adopts CogVAE.
H Reason for Using CogVAE without the Pre-trained Diffusion Model
@@
We conducted a comparative analysis of VAEs, as detailed in Section 5.3 and Appendix G, and found CogVAE [42] to perform well. Given that the VAE determines the upper limit of generation quality, we opted to use CogVAE for video encoding. Notably, CogVideoX [42] employs a novel DiT structure, where each layer’s latent space integrates both video and text condition information. This approach may complicate the design of geometry-related conditions. Furthermore, CogVideoX was not trained in a driving scenario. To eliminate these potential confounding factors, we decided to train the diffusion model from scratch using CogVAE. This strategy allows us to move beyond the constraints of pre-trained models, enabling more flexible modifications to the model architecture to achieve multi-view consistency and spatiotemporal encoding of geometry conditions.
@@
I Single Inference v.s. Rollout Inference
(a) Generation from Vista. It takes the first frame as input and generates the following (only supporting the front view).
@@
Figure III: Comparison between Rollout for Long Videos (Vista [12]) and Single Inference (our MagicDriveDiT). Although rollout can handle long videos, the quality is significantly degraded. In contrast, our extrapolation maintains high quality in long video generation.
@@
J More Visualization
As said in the “Note”, MagicDriveDiT is capable of generating $6\times 848\times 1600\times 241$ videos (20s at 12 fps). We include more generated samples in Figure IV-V. Please see the videos on our project page.
Figure IV: We show some frames from the generated $6\times 848\times 1600\times 241$ videos with the same scene configuration (i.e., boxes, maps, cameras, and ego trajectory) but under different weather conditions. Conditions are from the nuScenes validation set.
Figure V: We show some frames from the generated $6\times 848\times 1600\times 241$ videos with the same scene configuration (i.e., boxes, maps, cameras, and ego trajectory) but under different time-of-day conditions. Conditions are from the nuScenes validation set.